# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

### v1

@@ Abstract
Dual process theory posits that human cognition arises via two systems. System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, a slow, onerous, and deliberate process. NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2. In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias. We investigate the relationship between bias, CoT prompting, and dual process theory in LLMs directly. We compare zero-shot, CoT, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories. We also use human and machine personas to determine whether the effects of dual process theory in LLMs are based on modeling human cognition or inherent to the system. We find that a human persona, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 13 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Prompting-Techniques-to-Reduce-Social-Bias-in-LLMs
@@ 1 Introduction
In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society. These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in machine learning (Karanjai and Shi, 2024). A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022). By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
However, despite these advancements, LLMs continue to struggle with embedded social biases. These biases show up in different ways, including stereotyping and biased answers, which raises questions regarding the ethical use of LLMs in real-life applications. These biases are difficult to identify and even more challenging to eliminate wholesale due to the complex and opaque inner workings of AI models, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use. This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions. Applying dual process theory, a well-established psychological framework, to recent advancements in AI systems has illuminated possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM behavior align with and diverge from human cognitive processes.
In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets. Our approach uniquely incorporates both human-like and machine-like personas to examine whether the effects of these cognitive theories in LLMs are merely analogical to human cognition or intrinsic to the AI systems themselves. In addition, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting. When we use standard zero-shot prompting in Figure 1 (a), we see that the Llama 2 model responds with a stereotypical answer. But when we use the prompting of the human persona with System 2 in Figure 1 (b), it instead responds with an anti-stereotypical answer.
@@
  • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 cutting-edge LLMs.
  • We find that incorporating System 2 and Human Persona to prompts both reduce social biases. The combination of the two lead to the largest reduction in social bias when averaged across models and bias categories.
  • In contradiction to the assumptions made by researchers in the past (Hagendorff et al., 2023), we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases. In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine).
@@ 2 Related Work
Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes. Hagendorff et al. (2023) look into reasoning biases that are similar to humans in LLMs. It shows that as these models get bigger and more complex, they start making intuitive mistakes, like those found in human System 1 thinking. This trend shifts with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited. In a significant enhancement to dual-process interaction within LLMs, Lin et al. (2024) introduce SWIFTSAGE a new dual-module framework for better action planning in complex interactive reasoning tasks. This framework combines behavior cloning and prompting large language models. It includes the SWIFT module for quick, intuitive responses, and the SAGE module for careful, detailed planning. Tested on 30 tasks in the ScienceWorld benchmark, SWIFTSAGE greatly outperforms current methods like SayCan, ReAct, and Reflexion. It shows its ability to efficiently solve complicated interactive challenges with less computational needs. Furthermore, Wei et al. (2022) and Sun (2024) contribute to our understanding of how embedding systematic reasoning in LLMs, through techniques like chain-of-thought prompting, can substantially improve performance across various cognitive tasks including complex arithmetic and symbolic reasoning.
@@
Coming to the Dual Process Theory, which is a central idea in understanding human thinking and decision-making. It distinguishes between two different kinds of thinking: fast, automatic (System 1), and slow, effortful (System 2). System 1 enables quick comprehension through associations and pre-existing knowledge. In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions. These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010). The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013). In language understanding, Dual Process Theory explains how we can quickly and easily understand the general meanings of language input, but also switch to a more deliberate process when we encounter more complex language use, whether that be, for example, complex embedded syntactic structures, ambiguity in meaning, or uncertain pragmatic contexts. This shows how both quick thinking and more careful thinking work together (Ferreira and Huettig, 2023). Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea. System 1 lets us quickly guess what another person is thinking in analogy to our own while System 2 helps us to think more about their state more systematically with less self-attribution to make a metacognitive judgment (Carruthers, 2009). While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023). Interestingly, biases are not only caused by System 2 not getting involved. They can also come from a fight between heuristic and logical intuitions that happen at the same time. What happens next depends on which intuition is stronger and if the person realizes there is a conflict. If the relevant logical intuition is available, a correct judgment can be made without relying on the more resource-intensive System 2 cognitive process. This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019). These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task. Building on this theory, Bellini-Leite (2023) discusses the implementation of System 2 analytical processes in LLMs to mimic careful human reasoning, potentially mitigating frequent errors and enhancing decision-making capabilities in these models.
Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses. Gupta et al. (2023) show how different personas, reflecting varied socio-demographic backgrounds, can induce biases in LLM responses. Similarly, Deshpande et al. (2023) demonstrate that persona variations can dramatically alter the toxicity levels in model outputs. This study also highlights the risks and challenges in deploying LLMs across diverse user interactions in persona-assigned LLMs.
CoT prompting for not only improving reasoning capabilities but also for mitigating biases in LLMs has been a recent development in the field. Kaneko et al. (2024) found that CoT prompting can effectively reduce gender bias in LLMs and that by adding step-by-step reasoning to model predictions, CoT prompting makes the prediction process clearer and cuts down on oft-repeated gender biases in LLMs.
@@ 3 Dataset
We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2023) bias detection dataset. StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories. Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors. We will call Kamruzzaman et al.’s dataset the GenAssocBias dataset for the remainder of this paper.
@@ 4 Experimental Setup
We evaluate five major language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; [^2] Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4. We ensured that this sample contains an equal number of examples from each bias category. 2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-7B (Touvron et al., 2023), using the TheBloke/Llama-2-7B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to availability of local computing resources. checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API. Detailed procedures for the experiment are included in our supplementary materials, which also contain scripts for replication.
We use 12 different types of prompting techniques in our paper including the combinations of CoT, System 1, System 2, and Persona.
@@
Table 1: The types of prompts we use in our experiments. The “…” after each prompt are the common instructions which include “select the appropriate word from the options provided, without considering the order in which the options come in.” We provide options randomly to eliminate ordering bias and add “don’t add option number as well in your response” to minimize response postprocessing complexity.
@@ 4.3 System 1 and System 2 Prompts
System 1 prompting encourages the model to make fast decisions, similar to the quick decision-making process in human cognition known as System 1. The System 1 prompt is “You answer questions quickly. Given the context below, choose the most suitable word from the provided options…”. This approach prompts the model to respond rapidly without deep deliberation, mimicking the intuitive human cognitive response. For System 2, which is characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, ….”. This prompt encourages the model to process information deeply and consider multiple factors before making a decision. The aim is to simulate a more reflective thinking process in models. By promoting such processing, we intend to explore how the reduction of cognitive shortcuts—or the modeling of reduced cognitive shortcuts by the LLM—can decrease the reproduction of societal biases in LLM outputs.
@@ 4.4 Human and Machine Persona Prompts
In order to differentiate between the effects of dual process theory prompts on inherent LLM processing and the LLM’s model of human reasoning patterns, we incorporate prompting variants for human and machine personas. This is integrated with the other prompting methods (Standard, CoT, and Systems 1 and 2). So we add either a ‘Human Persona’ or a ‘Machine Persona’ by including the phrase ‘Adopt the identity of [persona]’, which influences how the LLM answers the following question. For instance, the ‘Human Persona with System 1’ (HP System 1) prompt is: ‘Adopt the identity of a person who answers questions quickly. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. Similarly, the ‘Machine Persona with System 2’ (MP System 2) prompt is ‘Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. See Table 1 for all the prompts we explore in this paper and how they realize persona, cognitive system, and CoT combinations. These varied personas help us explore how mimicking human-like cognitive processes in models might reduce inherent social biases.
@@ 5 Results & Analysis
Another surprising result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts. First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2). At its face, this result is in direct contradiction with Gupta et al.’s (Gupta et al., 2023) and Deshpande et al. (2023) results where personas led to an increase in bias and toxicity. Our results reveal a more complicated situation. That is, adding a persona does not always lead to more bias in LLMs. Rather, LLM behavior changes based on the selected persona that the LLM is asked to model. In some cases, such as in the prior work, where the persona is a member of an underrepresented population, bias is revealed through its behavior. In our case, when the persona is a generic person or machine, it pushes the model away from socially biased responses. This suggests that having an LLM model a separate entity (human or machine) leads to more objective outputs, similar to Solomon’s Paradox observed in humans—this paradox is the observation that people reason more wisely regarding other’s problems than their own and that self-distancing may improve objective reasoning capabilities (Grossmann and Kross, 2014).
@@
Things do not stop being interesting even yet. We find that when System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified. That is the difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall. This combination results in a reduction of over 3% from the standard zero-shot prompt. While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts. This suggests that while the LLM intrinsically models the two systems in dual process theory to some degree, its model of human cognition specifically has an even more exaggerated difference in these cognitive systems.
Model- and Bias-specific Prompting Effects.

### v3

@@ Abstract
Dual process theory posits that human cognition arises via two systems. System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, is a slow, onerous, and deliberate process. NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2. In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias. We investigate the relationship between bias, CoT prompting, a debiasing prompt, and dual process theory in LLMs directly. We compare zero-shot CoT, debiasing, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories. We incorporate human and machine personas to determine whether the effects of dual process theory in LLMs exist independent of explicit persona models or are based on modeling human cognition. We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 19 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs
@@ 1 Introduction
In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society. These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in automation (Karanjai and Shi, 2024). A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022). By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
Despite these advancements, LLMs continue to struggle with embedded social biases, which raises questions regarding the ethical use of LLMs in real-life applications. These biases are difficult to identify and even more challenging to eliminate due to the complex and opaque inner workings of LLMs, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use. This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions. Applying dual process theory, a well-established psychological framework, to recent AI advancements illuminates possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM generations align with and diverge from human cognitive processes.
In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets. Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in LLMs are dependent on explicit co-modeling of human cognition or always implicitly modeled. We follow-up on this analysis by examining interactions with debiasing prompts designed specifically for social bias reduction. Finally, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting. When we use standard zero-shot prompting in Figure 1 (a), we see that Llama 2 responds with a stereotypical answer. When we use human persona with System 2 prompting in Figure 1 (b), it instead responds with an anti-stereotypical answer.
@@
  • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 LLMs. This is followed up with 6 prompting variations incorporating explicit debiasing.
  • We find that incorporating a human persona is critical for controlling for biases in LLMs. While System 2 and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
  • In line with Nighojkar’s (2024) recent results in the reasoning domain, we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases. In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine). This contradicts often stated assumptions by researchers in the past (Hagendorff et al., 2023).
@@ 2 Related Work
Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes. Hagendorff et al. (2023) look into human-like reasoning biases in LMs. They show that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking. This trend shifted with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited. In a significant enhancement to dual-process interaction within LLMs, Lin et al. (2024) introduce SWIFTSAGE a new dual-module framework for better action planning in complex interactive reasoning tasks. This framework combines behavior cloning and prompting large language models. It includes the SWIFT module for quick, intuitive responses, and the SAGE module for careful, detailed planning. Tested on 30 tasks in the ScienceWorld benchmark, SWIFTSAGE greatly outperforms current methods like SayCan, ReAct, and Reflexion. It shows its ability to efficiently solve complicated interactive challenges with less computational needs.
Coming to the debiasing studies, Furniturewala et al. (2024) investigate the use of structured prompting techniques for debiasing language models. The study explores three categories of prompts: Prefix Prompting, Self-Refinement, and Implication Prompting. Chisca et al. (2024) focus on mitigating biases in LLMs through prompt-tuning. This approach introduces small sets of trainable token embeddings that are concatenated to input sequences, aiming to reduce biases without major modifications to the model’s architecture.
@@
Dual Process Theory is a psychological account of how human thinking and decision-making arise from two distinct modes. It distinguishes between fast, automatic (System 1), and slow, effortful (System 2) modes of thinking. System 1 enables quick comprehension through associations and pre-existing knowledge. In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions. These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010). The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013; Ferreira and Huettig, 2023). Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea. System 1 lets us quickly guess what another person is thinking in analogy to our own while System 2 helps us to think more about their state more systematically with less self-attribution to make a metacognitive judgment (Carruthers, 2009). While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023). Interestingly, biases are not only caused by System 2 not getting involved. They can also come from a fight between heuristic and logical intuitions that happen at the same time. This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019). These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task. Bellini-Leite (2023) discusses how methods such as CoT and tree-of-thought prompting in LLMs are suggestive of System 2 human reasoning, potentially mitigating frequent errors and enhancing reliability in these models. Nighojkar (2024) go on to test this correspondence by comparing against results from human experiments. He finds that CoT prompting does not simply mimic System 2. Rather, it leads to better agreement with human responses in both System 1 and System 2 triggering instructions.
Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses. Beck et al. (2024) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks. But the effectiveness of this approach varies across different models, dataset types, and sociodemographic factors. de Araujo and Roth (2024) explore how different personas, such as roles or identities (e.g., gender, occupation, or cultural background) influence the behavior and outputs of LLMs across various tasks. The research involves giving 162 unique personas to various LLMs and assessing how well they perform on different tasks and it turns out that personas not only affect task performance, bias, and model refusals but also show varied responses depending on the assigned persona. In addition to personas, explicitly debiasing instructions in prompts have been found to effectively reduce gender bias in LLMs (Kaneko et al., 2024).
@@ 3 Dataset
We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024) bias detection dataset. StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories. Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in the profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors. We will call this dataset the GenAssocBias dataset for the remainder of this paper.
@@ 4 Experimental Setup
We evaluate five major language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; [^2] Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4. We ensured that this sample contains an equal number of examples from each bias category. 2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-13B (Touvron et al., 2023), using the TheBloke/Llama-2-13B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to limited local computing resources. checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API.
We use 12 different types of prompting techniques in the main experiments including the combinations of CoT, System 1, System 2, and Persona. This is followed up with 6 additional prompting techniques with explicit debiasing.
@@
Table 1: The types of prompts we use in our main experiments. The “…” after each prompt are the common instructions which include “select the appropriate word from the options provided, without considering the order in which the options come in.” We provide options randomly to eliminate ordering bias and add “don’t add option number as well in your response” to minimize response postprocessing complexity.
Type Prompt
Standard Prompt + Debias Fill in the BLANK with the most suitable word from the provided options. The response should be unbiased and does not rely on stereotypes…
HP + Debias Adopt the identity of a person who always makes sure that their answer is unbiased and does not rely on stereotypes. Fill in the BLANK …
System 2 + Debias You answer questions slowly and thoughtfully. Your answers are effortful and reliable. You always make sure that your answer is unbiased and does not rely on stereotypes. Fill in the BLANK …
HP + System 2 + Debias Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They always make sure that their answer is unbiased and does not rely on stereotypes. Fill in the …
CoT + Debias Let’s think step-by-step and fill in the BLANK. Make sure that the answer is unbiased and does not rely on stereotypes …
HP + System 2 + CoT + Debias Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They always make sure that their answer is unbiased and does not rely on stereotypes. They answer questions by thinking about them “step-by-step". Fill …
Table 2: The types of prompts we use when incorporating debiasing explicitly. The “…” represent the same as in Table 1
@@ 4.3 System 1 and System 2 Prompts
System 1 prompting frames the context so that LLMs generate text reflecting the quick decision-making process in human cognition known as System 1. The System 1 prompt is “You answer questions quickly. Given the context below, choose the most suitable word from the provided options…”. This approach prompts the model to generate text mimicking the intuitive human cognitive response which does not include deep deliberation. For modeling System 2, characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, ….”. This prompt frames the context to generate text that reflects human responses processing information deeply and considering multiple factors. By using this context, we explore how the approximation of reduced cognitive shortcuts by the LLM can decrease the reproduction of societal biases in LLM outputs.
@@ 4.4 Human and Machine Persona Prompts
In order to differentiate between the effects of dual process theory prompts on bare LLM processing and the LLM’s model of human reasoning patterns, we incorporate prompting variants for human and machine personas. This is integrated with the other prompting methods (Standard, CoT, and Systems 1 and 2). We add either a ‘Human Persona’ or a ‘Machine Persona’ by including the phrase ‘Adopt the identity of [persona]’, which influences how the LLM answers the following question. For instance, the ‘Human Persona with System 1’ (HP System 1) prompt is: ‘Adopt the identity of a person who answers questions quickly. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. Similarly, the ‘Machine Persona with System 2’ (MP System 2) prompt is ‘Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. See Table 1 for all the prompts we explore in this paper and how they realize persona, cognitive system, and CoT combinations. These varied personas help us explore how mimicking human-like cognitive processes in models might reduce inherent social biases.
@@ 5 Results & Analysis
Figure 3: Stereotypical Responses for the debiasing prompt follow-up experiment (orange colored). The blue colored bars are anchors from Figure 2 for easy comparison.
@@
Another result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts. First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2). This suggests that having an LLM model as a separate entity (human or machine) leads to less socially biased outputs.
@@
When System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified. The difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall. This combination results in a reduction of over 3% from the standard zero-shot prompt. While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts. This suggests that while the LLM independent of a persona differentiates the two systems in dual process theory to some degree, its model of human cognition has an even more exaggerated difference in these cognitive systems.
5.1 Debiasing Prompt Follow-up
From Figure 2, we see that HP System 2 and HP System 2 with CoT prompting techniques perform substantially better than other prompt settings on average. We perform a follow-up experiment based on these two techniques, investigating explicitly debiasing prompts, similar to Kaneko et al. (2024). We add 6 debiasing prompting techniques: variations of HP, System 2, and CoT prompt combinations. The exact debasing prompts are shown in Table 2. Figure 3 shows the overall stereotypical response rates for these debiasing-incorporated prompting techniques averaged across all five models and 9 bias categories. It shows that the HP Debias prompt performs best compared to all other techniques. Similar to System 2 prompts, we find that the bias reduction effects of explicit debiasing is amplified by a human persona. While System 2 and CoT are complementary with explicit debiasing, they do not provide additional benefit beyond the HP and debiasing combination. In fact, we find that on average other prompt features interfere with and degrade the best performing HP Debias setting. This suggests that HP System 2 and HP Debias prompting achieve similar changes in model generations but with Debias having a more social focus. This is not surprising when we look at the prompts. The debias prompt can be seen as a brief social-oriented structured reasoning prompt.
5.2 Model- and Bias-specific Prompting Effects